During 2017, Campbell Manufacturing Company incurred $62,400,000 of research and development (R&D) costs to create a long-life battery to use in computers. In accordance with FASB standards, the entire R&D cost was recognized as an expense in 2017. Manufacturing costs (direct materials, direct labor, and overhead) are expected to be $62 per unit. Packaging, shipping, and sales commissions are expected to be $9 per unit. Campbell expects to sell 1,300,000 batteries before new research renders the battery design technologically obsolete. During 2017, Campbell made 433,000 batteries and sold 392,000 of them.

Required

a. Identify the upstream and downstream costs.

b. Determine the 2017 amount of cost of goods sold and the ending inventory balance that would appear on the financial statements that are prepared in accordance with GAAP.

c. Determine the sales price assuming that Campbell desires to earn a profit margin that is equal to 25 percent of the total cost of developing, making, and distributing the batteries.

d. Prepare a GAAP-based income statement for 2017. Use the sales price developed in Requirement c.
Answer #1

a) The upstream costs are costs incurred before the start of production and downstream costs are costs incurred with the start of production. Thus in the given question the research and development costs are upstream costs and manufacturing costs, packaging, shipping and sales commision are downstream costs.

Upstream Costs (Amount in $)
Research and Development costs 62,400,000
Downstream Costs
Manufacturing Cost (433,000 batteries*$62) 26,846,000
Packaging, Shipping and Selling Commission (392,000*$9) 3,528,000

b) Cost of goods sold = ($62+$9)*392,000 batteries =$ 27,832,000

Ending Inventory = $62*41000 batteries = $2,542,000

c) Total cost of developing, making and distributing the batteries per battery = $62 + $9 + ($62,400,000/1,300,000) = $62+$9+$48 = $119

Sales Price = $119 + ($119*25%) = $148.75

d)   Campbell Manufacturin Company

Income Statement (Amount in $)

Sales Revenue (392,000 batteries*$148.75) (A) 58,310,000
Cost of goods sold (B) 27,832,000
Gross margin (A-B) 30,478,000
Research and development 62,400,000
Selling Expenses 0
Net income (loss) (27,832,000 - 62,400,000) (31,922,000)